OpenStudy (anonymous):

Convert both 16.1 km/h and 880 ft/min to meters per second and compare

OpenStudy (tkhunny):

Multiply by one (1)! Over and over and over. Use whatever conversion factor you know. I'll do the first one. \(\dfrac{16.1\;km}{hr}\cdot\dfrac{1000\;m}{1\;km}\cdot\dfrac{1\;hr}{60\;min}\cdot\dfrac{1\;min}{60\;sec}\) - Check those units, multiply and divide all the numbers and see where it lands!

OpenStudy (anonymous):

@tkhunny thanks

OpenStudy (tkhunny):

Aw... You're not going to show us the other one?